If 12 men working 8hrs a day can do a piece of work in 15 days, how many hrs a day must 20 men work in order to do it in 8 day?
Answer by stanbon(75887)

If 12 men working 8hrs a day can do a piece of work in 15 days, how many hrs a day must 20 men work in order to do it in 8 day?
----
hrs is indirectly related to # of days and to # of men
-----
h = k/(d*m)
------
Solve for "k" using
"12 men working 8hrs a day can do a piece of work in 15 days"
8 = k/(15*12)
k = 1440
------
Equation:
h = 1440/(d*m)
how many hrs a day must 20 men work in order to do it in 8 day?
-----------------
h = 1440/(8*20)
---
hrs per day = 9
===================
Cheers,
Stan H.
====================